# Enhancing Naturalness in LLM-Generated Utterances through Disfluency Insertion

### v1

@@ Abstract
Disfluencies are a natural feature of spontaneous human speech but are typically absent from the outputs of Large Language Models (LLMs). This absence can diminish the perceived naturalness of synthesized speech, which is an important criteria when building conversational agents that aim to mimick human behaviours. We show how the insertion of disfluencies can alleviate this shortcoming. The proposed approach involves (1) fine-tuning an LLM with Low-Rank Adaptation (LoRA) to incorporate various types of disfluencies into LLM-generated utterances and (2) synthesizing those utterances using a text-to-speech model that supports the generation of speech phenomena such as disfluencies. We evaluated the quality of the generated speech across two metrics: intelligibility and perceived spontaneity. We demonstrate through a user study that the insertion of disfluencies significantly increase the perceived spontaneity of the generated speech. This increase came, however, along with a slight reduction in intelligibility.
@@
Syed Zohaib Hassan SimulaMet Oslo, Norway syed@simula.no                        Pierre Lison Norwegian Computing Center Oslo, Norway plison@nr.no                        Pål Halvorsen SimulaMet Oslo, Norway paalh@simula.no
1 Introduction
@@
Disfluencies are an intrinsic component of spontaneous speech, often manifesting as hesitations, fillers (’uh,’ ’like’), or repeated words and phrases. They play an important role in human speech communication and are a common occurrence across all types of non-scripted dialogues Lickley and Bard (1998); Yaruss and Quesal (2002). Previous studies have highlighted that disfluent speech is often seen as more spontaneous and natural Kampf (2022). Pauses and hesitations have also been shown to serve as markers of spontaneous thought, enhancing the perceived naturalness of conversation (Saryazdi et al., 2021). Incorporating disfluencies into the outputs of conversational agents may therefore improve their perceived "naturalness".
@@
This perception of naturalness is an important consideration when building conversational systems designed to emulate human behaviours in various settings. Such systems are often described as virtual avatars and are employed in applications such as educational tools, gaming, and healthcare support Chheang et al. (2024); Qin and Hui (2023); Yan and Alterovitz (2024). Previous studies have shown that factors like appearance, speech, lip-sync, and a strong sense of presence are important for user engagement and effectiveness with these avatars Hassan et al. (2022); Salehi et al. (2022), along with the quality of speech synthesis Mattheyses and Verhelst (2015). Avatars have also been employed for the purposes of training human professionals to handle sensitive dialogues, such as delivering bad news in healthcare settings (Andrade et al., 2010) or conducting investigative interviews with children who have experienced abuse Pompedda et al. (2015); Dalli (2021); Baugerud et al. (2021). As emotional stress tends to decrease speech fluency (Buchanan et al., 2014), disfluencies are highly frequent in those stressful and challenging dialogues, and an avatar developed for such training purposes should seek to reproduce such conversational behaviours.
@@
This paper presents an approach that artificially inserts disfluencies in non-disfluent utterances, and passes on the results to a speech synthesis model that is well-suited for the generation of spontaneous speech. Our contributions are as follows:
  • Fine-tuning an LLM with LoRA to introduce disfluencies in dialogue utterances.
  • A user study to evaluate the effectiveness of the generated disfluencies by converting the disfluent text into audio clips for assessment.
2 Background
Shriberg (1994) showed that disfluencies in speech are not random but follow specific patterns. Disfluencies can be categorized into typical and atypical forms. Typical disfluencies include hesitations, fillers like "um" and "uh" which often signal pauses for thought or plan future speech. Atypical disfluencies include conversational phenomena such as repetition, substitution, insertion, and speech errors Yaruss and Quesal (2002).
Only a few previous works have explored the generation of disfluencies in spoken utterances. To our knowledge, Marge et al. (2010) presented the first system attempting to artificially inserting disfluencies in the outputs of a dialogue system. Although their motivation is similar to ours, their insertion mechanism relied on a set of heuristic rules instead of a fine-tuned LLM. They also did not explicitly assess the intelligibility of the synthesized utterances as part of their study.
Yang et al. (2020) proposed a Planner-Generator model for generating natural and diverse disfluent texts, with the goal of creating synthetic data for training disfluency detection models. Their approach consists of two components: a Planner, which identifies optimal locations for inserting disfluent segments, and a Generator, which creates the corresponding disfluent phrases based on the Planner’s predictions. Their experiments demonstrated state-of-the-art performance for disfluency detection on the Switchboard corpus.
@@ 2 Background
Marie (2023) trained a disfluency generation model using the base version of the T5 model Raffel et al. (2020). Their base model was fine-tuned for few-shot learning on the Fisher corpus, using fluent-disfluent parallel data subsets of varying sizes and the full set of available utterances. The work presented in this paper extends this approach along several dimensions. In addition to using a larger, decoder-only model for the disfluency generation, we also provide a detailed human evaluation of how the insertion of disfluencies affects their perceived spontaneity and intelligibility.
3 Methods
@@
We outline below the dataset and fine-tuning method employed to insert disfluencies in non-disfluent utterances, as well as the text synthesis model converting those into speech.
3.1 Dataset
We used the Switchboard dataset including disfluency annotations Zayats et al. (2019). The Switchboard corpus was developed by transcribing telephonic conversations Godfrey et al. (1992), and disfluencies were initially hand-annotated in the Penn Treebank release Marcus et al. (1999). This corpus was subsequently refined by researchers to correct transcription errors Mantha (2000). Zayats et al. (2019) employed the cleaned-up, disfluency-tagged dataset, and re-annotated with BIO tagging to account for reparandum and correction spans. We also used the NXT Switchboard Corpus Calhoun et al. (2010), focusing on annotations about pitch contours, pauses, and other acoustic features. Our primary focus was on their annotations of silent pauses. These pauses often signal hesitation, planning, or shifts in dialogue.
@@ 3.1 Dataset
The data from Zayats et al. (2019) was used to extract instances of atypical disfluencies, while  Calhoun et al. (2010) provided instances of typical disfluencies. Table 1 shows an overview of both the training and test set. The rate of disfluency is defined as the ratio of disfluent tokens to the total number of tokens in an utterance.
@@ 3.2 Model Training
We fine-tuned the Llama-2-7b-chat-hf with LoRA Hu et al. (2021) for an efficient and targeted adaptation process. We chose Llama-2-7b to strike a balance between accuracy and computational efficiency. The fine-tuning setup included a maximum sequence length of 200 tokens, which represents a moderately long input that optimized the memory demands. For the LoRA-specific parameters, we set the LoRA rank (r) to 32 and the scaling factor (alpha) to 64, with a dropout rate of 0.1. The training was conducted with a batch size of 2 and gradient accumulation steps of 4, effectively simulating a larger batch size to stabilize training on limited computational resources. We adopt a learning rate of $2\times 10^{-4}$. The model was fine-tuned to generate disfluencies like repetition, substitution, insertions, and speech pauses, as well as typical disfluencies such as filled and silent pauses.
3.3 Text-to-speech
@@
Disfluencies are a speech phenomena. We use existing text-to-speech models to convert the disfluent utterances generated by the fine-tuned LLM into audio form. To select a TTS model for generating realistic disfluent speech, we conducted a pilot study in which we evaluated audio samples generated by three TTS models: Bark TTS model developed by Suno-AI Suno-AI (2023), Tortoise TTS Betker (2023) and OpenAI TTS OpenAI (2023). The evaluation focused on the prosody, intonation, and in particular the pronunciation of disfluencies. An important consideration was how the text-to-speech rendered false starts such as "Th- they" or "B- birthday", which are particularly frequent in Switchboard – and hence in the disfluent utterances produced by the fine-tuned LLM.
@@
Among the three models, Bark TTS was deemed most effective at synthesizing these disfluencies into speech. While all three TTS models did at times produce errors, such as buzzing sounds and prolong disruptive pauses in longer utterances, Bark TTS was found to perform better overall. To ensure the audio clips generated with Bark for the user study were of high quality, we regenerated samples when noisy artifacts such as buzzing sounds were detected in the audio outputs.
4 Automatic Evaluation
To assess the performance of our fine-tuned large language model (LLM) at inserting disfluencies, we first compare the disfluent utterances generated by the model with the actual utterances in the Switchboard test set, in which disfluencies and pauses are explicitly transcribed (Meteer et al., 1995). This comparison is conducted using both BLEU and BERTScore Zhang et al. (2019) metrics to evaluate lexical and semantic alignment. Table 2 summarizes the evaluation results for the fine-tuned LLM. The BLEU score suggests substantial lexical overlap between the generated outputs and reference transcriptions. The high BERTScores (precision, recall and $F_{1}$) also demonstrates the model ability to maintain a large degree of semantic consistency between the generated and reference texts. The average disfluency rate for generated disfluent text is 29.1%, which is slightly higher than found in training data, shown in Table 1.
Metric Score
BLEU Score 0.5524
BERTScore Precision 0.9287
BERTScore Recall 0.9370
BERTScore F1 0.9327
Table 2: Evaluation of the fine-tuned LLM on the task of replicating disfluencies in the Switchboard test set.
5 User Study
@@
To assess the effectiveness of our approach, we conducted a user evaluation study aimed at systematically understanding listeners’ perceptions of fluent versus disfluent audio conditions. These conditions were designed to simulate real-life conversational scenarios, allowing us to observe how variations in speech fluency impact listener experience.
5.1 Study Design
@@
In our user study, participants were asked to listen to 10 audio clips, five clips each for disfluent and fluent audios, of simulated conversations. Separate sets of conversation content were used to distinguish between the disfluent and fluent audio conditions. We used ChatGPT to generate ten fictive conversations based on real-life scenarios for human evaluation services, each with 5 turns of dialogues. We inserted disfluencies into five of these conversations using our fine-tuned language model, leaving the remaining five fluent. Both the fluent and disfluent conversations were converted into audio clips via the BARK TTS model. Conversations and detailed instructions can be found in the appendix. Participants were provided with the task to assess each audio recording based on two dimensions:
  1. 1.
    Intelligibility: How clear and comprehensible was the spoken conversation?
  2. 2.
@@
The study participants were asked to rate each clip on intelligibility and spontaneity on a 5-point Likert scale (from worst to best). The study involved 41 participants. They accessed the user study through an online questionnaire filled from their mobile phones (17), laptops (15), tablets (6), or desktop computers (3). 32 participants used headphones. Participation was voluntary, and each participant received a gift card as compensation.
5.2 Results
The average intelligibility ratings from the participants was $4.42\pm 0.79$ for fluent utterances and $4.23\pm 1.04$ for disfluent utterances. Spontaneity scored an average of $2.58\pm 1.58$ for fluent utterances and $2.91\pm 1.51$ for disfluent utterances. Figure 1 presents the mean and standard deviation scores for each metric under each condition. Fluent utterances demonstrated higher intelligibility and lower spontaneity compared to disfluent utterances.
Figure 1: Bar-plot (mean values and 95% conf. intervals) of user ratings on intelligibility and spontaneity.
The statistical significance of those results was assessed using both independent two-sample T-tests as well as mixed-effects models, with results shown in Table 3. For intelligibility, the $t$-statistic was $2.03$ ($p$ = $0.043$), and for spontaneity, the $t$-statistic was $-2.14$ ($p$ = $0.033$). Both $p$-values, below the $\alpha$ level of $0.05$, indicate statistically significant differences. As participants rated multiple clips in both conditions, we also derived a mixed effect analysis to account for within-subject variability and provide more accurate estimates of condition effects. For intelligibility, the mixed-effects model yielded a z-value of $2.809$ (p = $0.005$), confirming a statistically significant effect of the condition, consistent with t-test results. The group variance for intelligibility was $0.420$, indicating moderate variability across groups. For spontaneity, the model showed a z-value of $-2.615$ (p = $0.009$), again demonstrating a significant effect, with a group variance of $0.805$, reflecting greater between-group variability.
6 Discussion
Metric Test Statistic p-value GV
Intelligibility T-Test $t=2.03$ $0.043$
Mixed E. $z=2.809$ $0.005$ $0.420$
Spontaneity T-Test $t=-2.14$ $0.033$
Mixed E. $z=-2.615$ $0.009$ $0.805$
The two questions were sought to answer are whether intelligibility decrease when we add the disfluencies to the fluent spoken utterances and whether perceived spontaneity increase when we add the disfluencies to the fluent spoken utterances.
The main findings from this user study are that disfluencies significantly enhance the perceived spontaneity of speech, while leading to a slight reduction in intelligibility. The first effect is, however, stronger than the second, as evidenced by the results in Table 3. The statistical results obtained with the mixed-effect model also highlighted the variability within participants’ responses. Greater variability in spontaneity ratings suggests that perceptions of speech realism are more subjective and shaped by individual expectations of what constitutes a spontaneous interaction.
As shown in Table 2, the fine-tuned LLM effectively inserts disfluencies while preserving the content of the original utterance. High BERTScore and BLEU scores confirm its outputs align closely with reference transcriptions in the test set. Despite the sporadic nature of disfluencies in spontaneous speech, the model reliably mimics those in the Switchboard dataset. The disfluency rate (29.1%) slightly exceeds the training data (24.5%), indicating a tendency to overproduce disfluencies, which could be adjusted in future work.
7 Conclusion
This paper showed that (1) disfluencies can be automatically inserted into non-disfluent utterances using a fine-tuned LLM, and (2) this insertion enhances the perceived spontaneity of synthetic speech. More precisely, the ratings obtained from a user study revealed a trade-off: utterances in which disfluencies were inserted were deemed as more spontaneous by the participants, but also scored slightly lower on the intelligibility scale.
Those findings are particularly beneficial for the design of LLM-powered conversational avatars, such as those employed to train human professionals to handle "difficult" and stressful dialogues such as investigative interviews, in which disfluencies are particularly prevalent. Future work will focus on fine-tuning disfluency rates for specific scenarios, such as simulating stress or emotional speech in sensitive conversations.
8 Limitations
Although Bark TTS was the most effective TTS model for our study, it occasionally introduced artifacts like buzzing sounds or prolonged pauses, requiring manual intervention to regenerate samples in extreme cases. Additionally, the user study was limited to a small set of fictitious conversations generated by ChatGPT to simulate real-life scenarios. While representative, these scenarios may not fully capture the diversity of real-world conversational contexts where disfluencies occur. Both of these limitations could have influenced participants’ perceptions during the study, potentially affecting the results.
9 Ethical Considerations
The use of language models to introduce disfluencies into fluent utterances presents some ethical challenges. One significant concern is the potential for deception: by mimicking human behavior, these models create speech that "sounds more human," potentially tricking the listener into believing they are interacting with a human speaker when they are not. Although our overarching objective is to develop avatars for professional training purposes in which participants are well aware that they are interacting with a virtual agent, deploying these models in contexts where users are not informed of the artificial nature of their interlocutor could lead to ethical concerns about trust, consent, and manipulation.
Additionally, fine-tuning such models risks amplifying biases inherent in the training data, potentially associating disfluency patterns with specific linguistic or socio-demographic groups. There is also the risk that artificial disfluencies might be exploited to manipulate perceptions of a speaker’s credibility or emotional state, posing risks in sensitive domains such as legal or professional contexts.
To mitigate these risks, we advocate for clear disclosure policies in any deployment of such models that simulate human speech. Users should be informed whenever they are interacting with an artificial agent, ensuring transparency and preserving trust. Additionally, we are committed to the responsible release of our fine-tuned model and aggregated data, which will only occur after peer review and with accompanying guidelines to minimize potential misuse.
Appendix A Appendices
@@
Conversation No Conversation Data
@@
Conversation No Conversation Data

### v2

@@ Abstract
Disfluencies are a natural feature of spontaneous human speech but are typically absent from Large Language Models (LLMs) outputs. This absence can diminish the perceived naturalness of synthesized speech, which is an essential criterion when building conversational agents that aim to mimic human behaviors. We show how the insertion of synthetic disfluencies in the outputs can alleviate this shortcoming. The proposed approach involves (1) fine-tuning an LLM with Low-Rank Adaptation (LoRA) to incorporate various types of disfluencies into LLM-generated utterances and (2) synthesizing those utterances using a text-to-speech model that supports the generation of speech phenomena such as disfluencies. We assess the quality of the disfluency insertion through the use of automated metrics combined with a human study evaluating the disfluency insertion mechanism across two dimensions: intelligibility and perceived spontaneity. Token-level evaluation revealed that the model successfully learned patterns of disfluency placement but showed a tendency toward over-generation. Furthermore, the user study demonstrated that the insertion of disfluencies significantly increases the perceived spontaneity of the generated speech. This increase came, however, along with a slight reduction in intelligibility.
Keywords: disfluencies, spontaneous speech, fine-tuning, user study, virtual avatars
[…]
@@
Syed Zohaib Hassan1, Pierre Lison2, Pål Halvorsen1
1SimulaMet, Oslo, Norway
2Norwegian Computing Center, Oslo, Norway
{syed, paalh}@simula.no, plison@nr.no
Abstract content
1.  Introduction
@@
Disfluencies are an intrinsic component of spontaneous speech, often manifesting as hesitations, fillers (’uh,’ ’like’), or repeated words and phrases. They play an essential role in human speech communication and are common across all types of non-scripted dialogues Lickley and Bard (1998); Yaruss and Quesal (2002). Pauses and hesitations often serve as markers of spontaneous thought (Saryazdi et al., 2021), and previous studies have highlighted that disfluent speech is often perceived as more spontaneous and natural Kampf (2022).
@@
This perception of naturalness by human listeners is important when building conversational systems designed to emulate human behaviors in various settings. Such systems are often described as virtual avatars and are employed in applications such as educational tools, gaming, and healthcare support Chheang et al. (2024); Qin and Hui (2023); Yan and Alterovitz (2024). Previous studies have shown that factors like appearance, speech, lip-sync, and a strong sense of presence are essential for user engagement and effectiveness with these avatars Hassan et al. (2022); Salehi et al. (2022), along with the quality of speech synthesis Mattheyses and Verhelst (2015).
Avatars have also been employed to train human professionals to handle sensitive dialogues, such as delivering bad news in healthcare settings (Andrade et al., 2010) or conducting investigative interviews with children who have experienced abuse Pompedda et al. (2015); Dalli (2021); Baugerud et al. (2021). As emotional stress tends to decrease speech fluency (Buchanan et al., 2014), disfluencies are highly frequent in those stressful and challenging dialogues, and an avatar developed for such training purposes should seek to reproduce such conversational behaviors.
@@
This paper presents a new method for achieving this objective. More specifically, the proposed approach artificially inserts disfluencies in non-disfluent utterances and passes on the results to a speech synthesis model that is well-suited for generating spontaneous speech. The disfluency insertion is performed by an LLM specifically fine-tuned for this task using LoRA on the basis of the Switchboard corpus (Godfrey et al., 1992).
This paper presents three contributions:
  • The fine-tuning of an open-weight LLM to introduce disfluencies in dialogue utterances, which are then used as input to a text-to-speech model.
  • The evaluation of the quality of the disfluency insertion through reference-based metrics.
  • A user study to evaluate the naturalness and intelligibility of the disfluent speech generated by such a process.
Section 2 provides a brief overview over existing work on disfluencies and disfluency generation. Section 3 presents the approach developed to automatically insert disfluencies in non-disfluent utterances. This approach is then evaluated by comparing the resulting utterances to the ones found in Switchboard (Section 4) and by conducting a user study (Section 5). Finally, Section 6 discusses those empirical results, and Section 7 concludes the paper.
2.  Background
Shriberg (1994) demonstrated that speech disfluencies are not random but follow specific patterns. They can be categorized as typical (e.g., hesitations, fillers like "um" and "uh") or atypical (e.g., repetitions, substitutions, insertions, and speech errors Yaruss and Quesal (2002)). Table 1 shows examples of different types of disfluencies commonly observed in spontaneous speech.
Disfluency type Example
Repetition he he didn’t call me
Substitution she went to he went to the store
Filled Pause I saw uh the movie yesterday
Insertion well I was thinking of going there
Speech Error I remember your bir- birthday.
Table 1: Examples of speech disfluencies. Strike-through formatted words represent the parts that should be edited out to obtain a fluent utterance, and correspond to the reparandum and interregnun in Shriberg (1994)’s terminology.
A few previous works have explored the generation of disfluencies in spoken utterances. To our knowledge, Marge et al. (2010) presented the first system to insert disfluencies in dialogue system outputs, relying on heuristic rules instead of a fine-tuned LLM. Passali et al. (2022) introduced large-scale artificial disfluency (LARD), which generates disfluencies like repetitions, replacements, and restarts using a synthetic approach that selects random positions within fluent text sequences, modifying them algorithmically to create disfluent variants.
Yang et al. (2020) introduced a Planner-Generator model to generate diverse disfluent texts for training disfluency detection models. The Planner identifies optimal disfluency insertion points, while the Generator produces corresponding phrases.
@@ 2.  Background
Marie (2023) fine-tuned the base T5 model Raffel et al. (2020) for few-shot learning on the Fisher corpus, leveraging fluent-disfluent parallel subsets of varying sizes and the complete set of utterances. This paper builds on their approach and extends it along several dimensions. In addition to using a larger, decoder-only model for the disfluency generation, we also provide a detailed human evaluation of how the insertion of disfluencies affects their perceived spontaneity and intelligibility.
3.  Methods
@@
We now describe the fine-tuning approach employed to insert disfluencies in non-disfluent utterances, as well as the text synthesis model converting those into speech.
3.1.  Dataset
We used the Switchboard dataset with disfluency annotations by Zayats et al. (2019). The dataset was based on the original Switchboard corpus Godfrey et al. (1992), previously annotated by Marcus et al. (1999) and subsequently cleaned and re-annotated using BIO tagging to accurately capture reparandum and correction spans. We also used the NXT Switchboard Corpus Calhoun et al. (2010) to enrich the aforementioned data with explicit annotations of silent pauses, which often signal hesitation, planning, or shifts in dialogue.
@@ 3.1.  Dataset
We represent the training data as pairs of utterances $(u_{f},u_{d})$ where $u_{d}$ is the transcription of the original utterance (thus including disfluencies) in the Switchboard corpus, and $u_{f}$ corresponds to its non-disfluent equivalent after the removal of all disfluencies from the transcription. The data from Zayats et al. (2019) was used to extract instances of atypical disfluencies, while Calhoun et al. (2010) provided instances of typical disfluencies. Table 2 shows an overview of both the training and test sets. The rate of disfluency is defined as the ratio of disfluent tokens to the total number of tokens in an utterance.
@@ 3.2.  Model Training
We fine-tuned Llama-2-7b-chat-hf Touvron et al. (2023) and Flan-T5-Large  Chung et al. (2024) with LoRA Hu et al. (2021) to generate disfluent utterances from fluent inputs. The fine-tuning setup included a maximum sequence length of 200 tokens to balance accuracy with computational efficiency. We selected LoRA for its computational efficiency and parameter-effective approach, which updates only a small subset of rank-specific parameters rather than the entire model as in supervised fine-tuning (SFT). This minimized over-fitting risks while enabling targeted adaptations. To balance generalization and efficiency, we set LoRA rank (r) to 32, scaling factor (alpha) to 64, and dropout rate to 0.1. Training employed a batch size of 2 with gradient accumulation steps of 4, simulating larger batches on limited resources, and a learning rate of $2\times 10^{-4}$. The model was fine-tuned for two specific tasks: generating atypical disfluencies (repetition, substitution, insertions, speech pauses) and typical disfluencies (filled and silent pauses).
$3.$  Text-to-speech
@@
Disfluencies are a speech phenomenon. We use existing text-to-speech models to convert the disfluent utterances generated by the fine-tuned LLM into audio form. To select a TTS model for generating realistic disfluent speech, we conducted a pilot study in which we evaluated audio samples generated by three TTS models: Bark TTS model developed by Suno-AI Suno-AI (2023b), Tortoise TTS Betker (2023) and OpenAI TTS OpenAI (2023). The evaluation focused on the prosody, intonation, and, in particular, the pronunciation of disfluencies. An important consideration was how the text-to-speech rendered false starts such as "Th- they" or "B- birthday", which are particularly frequent in Switchboard – and hence in the disfluent utterances produced by the fine-tuned LLM.
@@
Among the three models, Bark TTS was deemed the most effective at synthesizing these disfluencies into speech. While all three TTS models sometimes produced errors, such as buzzing sounds and prolonged disruptive pauses in longer utterances, we found that Bark TTS provided the most natural-sounding speech when applied to our disfluent transcriptions. To ensure the audio clips generated with Bark for the user study were high quality, we regenerated samples when noisy artifacts, such as buzzing sounds, were detected in the audio outputs.
4.  Evaluation
To assess the performance of our disfluency insertion approach, we start with automatic evaluations of the fine-tuned models. More precisely, we examine the extent to which those models replicate disfluency patterns found in natural speech, by comparing overlaps between real disfluent utterances and their synthetically generated counterparts.
4.1.  Reference-based Metrics
We first compare the disfluent utterances generated by each fine-tuned model with the actual utterances in the Switchboard test set. We used BLEU and BERTScore as metrics: while BLEU Papineni et al. (2001) measures the lexical overlap between the generated outputs and the reference transcriptions, the BERTScores Zhang et al. (2019) indicates the model’s ability to maintain semantic consistency between the generated and reference texts. Table 3 presents the results. The llama-2-7B model achieved a higher BLEU score, indicating better lexical overlap with reference transcriptions. Conversely, flan-t5 showed stronger semantic precision but lower recall, resulting in a slightly higher overall $F_{1}$ score.
Metric llama-2 flan-t5
BLEU Score 0.5524 0.4440
BERTScore Precision 0.9287 0.9624
BERTScore Recall 0.9370 0.9168
BERTScore F1 0.9327 0.9386
Table 3: Performance of two fine-tuned LLMs in reproducing the disfluency patterns found in the Switchboard test set, measured using BLEU and BERTScore.
As llama-2-7b produced disfluent utterances that exhibited a higher overlap (based on BLEU) with human-produced disfluent utterances, we conducted a token-level evaluation for disfluency placement exclusively on this model to understand not only whether it generates appropriate disfluencies, but also where it places them relative to human-produced disfluencies, as explained below.
4.2.  Disfluency placement
Metric Value 95% CI
Precision 52.2% [51.4, 53.0]
Recall 85.8% [85.0, 86.6]
F1-Score 61.2% [61.0, 61.9]
Accuracy 81.6% [81.4, 81.8]
Jaccard Similarity 30.4% [29.8, 31.0]
AUC Score 0.603
Table 4: Token-level overlap between human-produced disfluencies and disfluencies synthetically generated with our model (NN=3,610 utterances, for a total of 96,573 tokens).
To determine whether the fine-tuned LLM (llama-2-7b) mimics disfluencies in the same locations and types as humans, we then compared the relative positions of disfluencies in human and LLM-generated utterances using standard token-level evaluation metrics (precision, recall, F1 score, and AUC score). We manually annotated 3,610 fluent-disfluent sentence pairs from our test set and utterances with LLM-generated disfluencies using a rule-based approach, marking each token position as either disfluent (D tag) or fluent (O tag) in both the human-produced disfluent utterances and the model-generated outputs. Tokens were marked as disfluent if they were filler words (e.g., "uh", "um"), cut-off words (ending with "-"), absent from the original fluent text, or repetitions of previous words; otherwise, they were tagged as fluent.
Table 4 presents the overall results of this comparison for token-level disfluency matching. Compared to the manually annotated disfluencies, the utterances with synthetic disfluencies generated by our approach demonstrated high recall but moderate precision, revealing a tendency towards over-generation. The Jaccard similarity coefficient averaged 30.4% (±27.5%), indicating moderate overlap between model predictions and human annotations.
4.3.  Type-Specific Accuracy
The overlaps between human-produced and synthetic disfluencies varied significantly between disfluency types. Table 5 factors this variation across typical disfluency insertions (SIL) and atypical disfluencies (DISFL). Typical disfluency insertions showed higher recall but lower precision than atypical disfluencies. The fine-tuned disfluency generation model generated substantially more typical disfluencies than humans ($12.9$ vs $5.0$ per sample on average), while atypical disfluencies showed more controlled generation patterns ($5.0$ vs $4.2$ per sample).
Metric SIL DISFL Overall
No. Samples 1,094 2,516 3,610
Precision (%) 49.5 53.4 52.2
Recall (%) 93.4 82.5 85.8
F1-Score (%) 61.6 61.1 61.2
Avg Disfl/Sample 12.9 5.0 7.4
Human Baseline 4.2 Avg Disfl/Sample
Table 5: Token-level overlap factored by disfluency type: silence insertions (SIL) or structural disfluencies (DISFL).
5.  User Study
@@
To assess how the insertion of disfluencies in utterances affected their perception by human listeners, we conducted a user evaluation study comparing fluent and disfluent speech conditions on two criterias: intelligibility and perceived spontaneity. We designed these conditions to simulate real-life conversational scenarios, allowing us to observe how variations in speech fluency affect the listener’s experience.
5.1.  Study Design
@@
In our user study, we asked participants to listen to 10 audio clips of simulated conversations, five clips each for disfluent and fluent audio. We used separate sets of conversation content to distinguish between the disfluent and fluent audio conditions. We used ChatGPT to generate ten fictive conversations based on real-life scenarios for human evaluation services, each with five turns of dialogue. We inserted disfluencies into five conversations using our fine-tuned language model, leaving the remaining five fluent. We converted the fluent and disfluent conversations into audio clips using the BARK TTS model. The instructions given to the participants are shown in Figure 1, along with the transcriptions of the (respectively fluent and disfluent) dialogues to assess in Tables 6 and 7.
Participants were requested to assess each audio recording based on two criteria:
  • Intelligibility: How clear and comprehensible was the spoken conversation.
@@
Conversation Conversation Data
@@
Conversation Conversation Data